The sum of present age of two brothers 36 years. After four years their age will be 5:6. The age of the older brother ?
Answer by lwsshak3(11628)

The sum of present age of two brothers 36 years. After four years their age will be 5:6. The age of the older brother ?
***
let x=present age of younger brother
36-x=present age of older brother
After 4 yrs:
x+4=age of younger brother
36-x+4=40-x=age of older brother
age of younger brother/age of older brother=5/6
(x+4)/(40-x)=5/6
6x+24=200-5x
11x=176
x=16
36-16=20
present age of older brother=20